OpenStudy (anonymous):

what is the answer to 5x = 2 over 3 ? HELP ME PLEASE!

OpenStudy (precal):

is this your problem?

OpenStudy (anonymous):

Divide 2/3 by 5 to get X alone

OpenStudy (anonymous):

yeah thats my problem

OpenStudy (anonymous):

how do i divide 2 / 3 by 5?

OpenStudy (johnweldon1993):

or it might be easier to think of it in terms of multiplying each side by 1/5 \[\frac{ 1 }{ 5 } 5x = \frac{ 2 }{ 3 }\frac{ 1 }{ 5 }\] \[x = \frac{ 2 }{ 3 }\frac{ 1 }{ 5 }\] so what is 2/3 times 1/5?

OpenStudy (anonymous):

2/ 15?

OpenStudy (anonymous):

Yep!

OpenStudy (johnweldon1993):

Perfect

OpenStudy (anonymous):

can it be simplified ?

OpenStudy (johnweldon1993):

can it? does any number go into both 2 and 15 evenly?

OpenStudy (anonymous):

no

OpenStudy (johnweldon1993):

Then that is your final answer :)

OpenStudy (anonymous):

how about 2k = 5/6? im only 12 and they gave me way too hard math work. :(

OpenStudy (johnweldon1993):

don't worry about it. Well remember how we just multiplied both sides by 1/5 to cancel out 5x? to cancel out 2k...what do we multiply by?

OpenStudy (anonymous):

1/2 times 5/6 ?

OpenStudy (johnweldon1993):

perfect...which comes out to?

OpenStudy (anonymous):

5/ 12?

OpenStudy (johnweldon1993):

and correct again :) and that cannot be simplified correct?

OpenStudy (anonymous):

correct?

OpenStudy (johnweldon1993):

...correct lol...so 5/12 is your answer!

OpenStudy (anonymous):

thank you!!!! youre sooooooo awesome!!! you should be a math teacher!

OpenStudy (anonymous):

I'm going to be one :)

OpenStudy (johnweldon1993):

I'm a tutor actually lol...and that's awesome @FutureMathProfessor good luck!!

OpenStudy (anonymous):

Me too in college xD

OpenStudy (johnweldon1993):

as am I lol....do it as part of a work-study thing lol